"You are traveling 100 miles to get to your destination. For the first half of the trip you go 65 mph, and the second half you go 75 mph. What is your average speed?"
This was a question in an a team practice and the answer was supposedly 69.6, even though we all thought the answer was 70. What is the real answer, and if it is 69.6, how?

Answer by TimothyLamb(4379)

s = d / t
t = d / s
---
a = (100/2)/65
a = 50/65
---
b = (100/2)/75
b = 50/75
---
total time:
t = a + b
t = 50/65 + 50/75
t = 50(75)/(65)75 + 50(65)/(65)75
t = 1.4358974 hours
---
average speed:
s = d / t
s = 100 / 1.4358974
s = 69.642857 mph
---
note that:
(65 + 75)/2 = 70 mph is NOT the correct answer
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php